Solve for x : √x + 1 - √x – 1 = 1.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. √x + 1 = 1 + √ x – 1
Squaring both sides , we get
x + 1 = 1 + x – 1 + 2 √x – 1 ⇒ 1 = 2√x – 1.
⇒ 1 = 4(x – 1)
⇒ x = 5/4